# Aristotle’s Second-Order Logic

### v11

@@ Introduction
and call the resulting system ASOL. It seems very plausible that AML(see [17]) can be embedded into ASOL and all the corresponding theorems be derived.
In this section we have provided evidence that ASOL is a very natural candidate to express the content of the Analytics and its metalogical inference and use of counterexamples and have given examples of reasonings which can be seen very naturally as deploying both first and second-order versions of the introduction and elimination rules for existential and universal quantifiers as well as definitions and propositions in which these two types of quantifiers occur in a non-trivial way over (second-order) binary predicates.
Natural Deduction in Aristotle and Galen
In this section we focus specifically and in more detail on evidence for Aristotle’s conscious understanding and deployment of the first-order natural deduction quantifier rules and supplement our argument with the valuable evidence provided by Galen. From what we said in the previous section Aristotle’s deployment of the second-order quantifier rules $\forall^{2}$ and $\forall^{2}$ is evident and the same goes for $\exists^{2}$, for instance in his frequent use of examples with concrete terms. Our discussion in the next two sections about the Topics provides further evidence. The use of $\exists^{2}$ still needs to be investigated.
We argue that in Aristotle (and Galen) there is evidence for i) the conscious (philosophically reflected) use of the first-order introduction and elimination rules for $\forall$ and ii) that these rules were applied to multiple universal quantification over non-necessarily monadic predicates.
[…]
In the Organon and other works Aristotle presents us with evidence of complex and accurate propositional reasoning. For instance in how constructive and refutation topics are inferred from each other, as will be illustrated in the following sections. This fact cannot be explained by mere natural language competency, mathematical training or a spontaneous manifestation of Aristotle’s specific intellectual gifts. Rather there must have been a significant continuity between Aristotle’s or his disciple Theophrastus’ (who wrote more logical treatises than his teacher) and Chrysippus’ patently sophisticated formal system of propositional logic[6].
In the Analytics universally quantified sentences are considered which include non-monadic predicates. Such was the typical form of an axiom or principle. Consider the discussion of the axioms that need to be applied in the geometric proof in the Prior Analytics 41b13–22. One axiom mentioned can be written directly in ASOL as
$\forall xyzw(x=y\enspace\&\enspace z=w\rightarrow x-z=y-w)$
[…]
$\forall abc(Double\,ab\enspace\&Double\,bc\rightarrow Quadruple\,ac)$
[…]
We now argue i) and ii) for the case of $\forall$-introduction. Consider the remarkable passage from chapter 3 of Book II of the Topics (tr. W.A. Pickard):
Whereas in establishing a statement we ought to secure a preliminary admission that if it belongs in any case whatever, it belongs universally, supposing this claim to be a plausible one. For it is not enough to discuss a single instance in order to show that an attribute belongs universally; e.g. to argue that if the soul of man be immortal, then every soul is immortal, so that a previous admission must be secured that if any soul whatever be immortal, then every soul is immortal. This is not to be done in every case, but only whenever we are not easily able to quote any single argument applying to all cases in common euporômen koinon epi pantôn hena logon eipein , as (e.g.) the geometrician can argue that the triangle has its angles equal to two right angles.
To explain this passage we make use of the ASOL interpretation of the Topics developed in the next section. What Aristotle is first saying (in the intensional reading) is that in general we do not have ’induction’ from the species to the genus. If (essential) accident (or difference) $C$ applies to species $A$ of $B$ it does not follow that accident $C$ applies to $B$. That is, from $\gamma^{\star}AC\&\sigma AC$ it does not follow that $\sigma BC$. Thus if the human soul is immortal it does not follow that the soul is immortal (extensionally: every soul is immortal). Aristotle says that we have to assume not only $\sigma AC$ (and $\gamma^{\star}AB$) but also that $\gamma^{\star}AC\&\sigma AC\rightarrow\sigma BC$. Aristotle contrasts this situation to another situation in which euporômen koinon epi pantôn hena logon eipein - in which we are able to deduce (in a uniform way) a universal conclusion through a single argument, as the geometrician reasons about property of a triangle. We interpret this to mean, as seems natural, that this ’single argument’ is one involving a variable $X$ over which the universal conclusion is quantified. Thus the geometrician may start with ’Let $X$ be a triangle’ and arrive at the conclusion $X$ has its angles equal to two right angles and thereby view this as a proof that all triangles have their angles equal to two right angles. Thus we are lead to interpret Aristotle’s koinon epi pantôn hena logon eipein as a statement of the rule $\forall^{1}$ (and $\forall^{2}$).
It is also interesting to compare this to Proclus’ commentary of the first book of Euclid [19], p.162 which may be also interpreted as expressing the rule $\forall^{1}$:
[…]
The previous observations lead us to conclude that ancient logic had at least the full expressive and inferential power of open theories. Open theories (sometimes called ”quantifier free” theories) are far from trivial as they include the theories of important algebraic structures such as real closed fields. A considerable portion of arithmetic can be formalized in the quantifier-free fragment of primitive recursive arithmetic [24]. It is also interesting to note (in our ASOL context) that in second-order logic we can define both first and second-order existential quantifiers in terms of universal quantifiers alone [15], p.67.
Similarly as for universal quantification we will attempt to show that there was a conscious use of the introduction and elimination rules for first-order existential quantification and specifically for such quantifiers applied to non-monadic predicates.
Consider again the proof of the conversion of the universal negation in the Prior Analytics 25a14-18 discussed in section 2 (tr. Hardie and Gaye):
First then take a universal negative with the terms A and B. If no B is A, neither can any A be B. For if some A (say C) were B, it would not be true that no B is A; for C is a B. But if every B is A then some A is B. For if no A were B, then no B could be A. But we assumed that every B is A
Aristotle’s argument (expressed in a rather terse and laconic style) can be read as follows: If A is not said of any B then neither is B said of any A. For assume that B is said of some A (Aristotle takes this immediately to be equivalent to it not being the case that B is not said of any A). Let this some A be $c$. Then $c$ is a $A$ which is also a $B$. But this contradicts the initial hypothesis that A is not said of any B. Hence by reductio ad absurdum we conclude that it is not the case that B is said of some A, that is to say, B is not said of any A.
The proof (which makes use of $\bot_{c}$) can be interpreted directly as follows in linearized natural deduction, in which we have condensed some logical steps and assumed basic Aristotelian premises in the Prior Analytics:
1. $\forall x(Bx\rightarrow\sim Ax)$ Hyp
2. $\sim(\forall x(Ax\rightarrow\sim Bx))$ Hyp
3. $\exists xAx\&Bx$ 2
4. $Ac\&Bc$ Hyp
5. $\bot$ 1,4
6. $\bot$ $\exists^{1}E$ 3-5
7. $\forall x(Ax\rightarrow\sim Bx)$ $\bot_{c}$ 2, 6
8. $\forall x(Bx\rightarrow\sim Ax)\rightarrow\forall x(Ax\rightarrow\sim Bx)$ $\enspace\rightarrow I$ 1,7
In 6 hypothesis 4 is discharged. Besides $\bot_{c}$ (which is common and consciously employed in Aristotle) the key ingredient is $\exists^{1}E$. The evidence points to $c$ being used in Aristotle’s proof exactly as $c$ in the natural deduction rule $\exists^{1}E$.
In 28a24-27 of the Prior Analytics Aristotle gives two proofs a certain syllogism in the third figure, the second one which employs a process called tô(i) ekthesthai poiein, byekthesis:
Recall that for Aristotle particular predication has existential import so we assume $\exists xSx$ as an axiom (NEmp). Aristotle’s second proof of this third figure figure (by exposition or ekthesis) can be read in detail as follows. Assume that all S is P and all S is Q. We have that there exists an S. Let $n$ be such an $S$. Then by the hypotheses we conclude that $n$ is $P$ and also that $n$ is $Q$. Thus some $P$ is $Q$.
The proof can be read as follows in linearized natural deduction:
1. $\forall x(Sx\rightarrow Px)\enspace\&\enspace\forall x(Sx\rightarrow Qx)$ Hyp
2. $\exists xSx$ Ax
3. $Sn$ Hyp
4. $Sn\supset Pn\enspace\&\enspace Sn\rightarrow Qn$ $\forall^{1}E$ 1
[…]
6. $\exists xPx\&Qx$ $\exists^{1}I$ 5
7. $\exists xPx\&Qx$ $\exists^{1}E$ 2-6
8. $\forall x(Sx\rightarrow Px)\enspace\&\enspace\forall x(Sx\rightarrow Qx)\rightarrow\exists xPx\&Qx$ $\enspace\rightarrow I$, 1, 7.
In 7 we discharged hypothesis 3. Again we have similar evidence for the deployment of the rule $\exists^{1}$ but this time in conjunction with $\exists^{1}$.
We propose that for Aristotle proof by ekthesis involved (essentially) the use of $\exists^{1}E$ in conjunction with other rules, perhaps mainly $\exists^{1}$. However a case can be made that Aristotle deployed reasoning in cases in which $\exists^{I}$ would be deployed which amounts to considering Skolemizations or Choice Functions of the corresponding formulas.
Proclus considers constructions as a special type of proof. Unlike in modern ’classical’ mathematics Euclid does to revel in non-constructive existence theorems. A modern formalization of the first proposition of Book I would take the form $\forall x\exists yTxy$. This becomes through Skolemization
$\exists f\forall xT(x,fy)$
which is shown constructively by effectively producing such a $f$. So what is explicitly proven in Euclid is $\forall xT(x,fy)$, that for every segment $\overline{AB}$ the construction $f(\overline{AB})$ produces an equilateral triangle. It would seem that Proclus offers evidence that the ancients anticipated modern intuitionistic or constructivism interpretations of the existential quantifier666see[11] for further discussion on the constructivist interpretation of the existential quantification in Euclid. In appendix 3 we sketch how one might go about formalizing Euclid in ASOL in a way which seems natural and which also seems to fit within an intuitionistic or constructivist framework. in situations of multiple generality. It could be that Skolemization or Choice Functions might be quite an important aspect of the logical treatment of propositions of the form $\forall X\exists Y\phi(X,Y)$ (here $X$ and $Y$ are either individual variables or predicate term variables) in Aristotle and other ancient authors, as an alternative to the deployment of natural deduction rule $\exists^{1}E$. For $\forall X\exists Y\phi(X,Y)$ we would have a choice function $j(X)$, where $\phi(X,j(X))$. This is like an indefinite dependent version $\epsilon$ of the Peano operator and $j(X)$ can be read as a $Y$ such that $\phi(X,Y)$. We can write $j(X)=\epsilon Y\phi(X,Y)$. In this way we interpret this constructions: let $a$ be a $Y$ such that $\phi(X,Y)$ or in more natural language form: let $a$ be a $\phi$ of $X$.
And we can consider this to include the case of $\exists X\phi(X)$ where we deploy the indefinite version of the Peano operator (which corresponds to a definite article): $a=\epsilon X\phi(X)$, let $a$ be a $\phi$.
Consider proposition 2777Given two unequal magnitudes, it is possible to find two unequal straight lines such that the greater straight line has to the less a ratio less than the greater magnitude has to the less (Heath tr.). of Archimedes’s On the Sphere and Cylinder (Book I). Archimedes assumes that given quantities $A<B$ there is natural number $n$ such that $n\bullet A>B$ where $n\bullet A$ means the quantity obtained by adding $A$ $n$ times to itself. Thus this axiom reads $\forall A,B\exists n(A<B\rightarrow(n\bullet A>B))$. The way Archimedes uses the axiom in the proof [^8] The relevant part of the proof reads (Heath tr.): Then, if $CA$ is added to itself a sufficient number of times, the sum will exceed $D$. Let $AF$ be this sum. This is equivalent to: let $AF$ be a multiple of $CA$ larger than $D$. is far more suggestive of the deployment of a Choice Function $j$ for the axiom than $\exists^{1}E$. Archimedes takes directly a $C=j(A,B,n)\bullet A$ where $j$ is a Choice Function for the axiom.
We could read the ekthesis in 25a14-18 examined above as follows (and similarly for 28a24-27):
1. $\forall x(Bx\rightarrow\sim Ax)$ Hyp
2. $\sim(\forall x(Ax\rightarrow\sim Bx))$ Hyp
3. $\exists xAx\&Bx$ 2
4. $c=\epsilon x(Ax\&Bx)$ 3
5. $Ac\&Bc$ by the expected axiom for $\epsilon$ and 3
6. $\bot$ 1,5
7. $\forall x(Ax\rightarrow\sim Bx)$ $\bot_{c}$ 2, 6
8. $\forall x(Bx\rightarrow\sim Ax)\rightarrow\forall x(Ax\rightarrow\sim Bx)$ $\enspace\rightarrow I$ 1,7
[…]
Sentences of the form $\forall x\exists yRxy$ abound in Aristotle. The (modern) concept of function corresponds to the case $\forall x\exists^{1}yRxy$ and is found embedded in Aristotle’s Physics. It appears in context of proofs in expressions in the grammatical form ’let $a$ be the $y$ such that $Rxy$’ (i.e. the time taken of a given motion) wherein it is understood that $\forall x\exists^{1}yRxy$ is assumed to hold. More generally we meet expressions of the form $\forall x,y\exists^{1}zMxyz$ and for a given $b$ and $m$ we find: let $t$ be the $z$ such that $Mbmz$ (which corresponds to the $\epsilon$ construction discussed above). For instance: let $t$ be the time taken by body $b$ in motion $m$. The proofs can be interpreted as deploying $\exists^{1}$, for instance using $\exists yRxy$ and hypothesis $Rxa$. In the Physics there is an implied functional correspondence, for a given body, between all the continua of motion, space, time and magnitude. A division of one of these corresponds to a division of the others. The principle manner in which such suppositions are used would seem to correspond to the deployment of $\exists^{1}$ together with other rules - which is our proposed interpretation of ekthesis. We propose that ekthesis may also be interpreted in terms of the $\epsilon$ operator (an interpretation we will develop in a future work). This technical term can be read as a sign that for Aristotle the deployment of the natural deduction rules for the existential quantifier was conscious and reflected. In this ’functional’ case of unique existential quantification we have unique choice, reflected in the corresponding definite article in the grammatical construction: let $y$ be the $x$ such that $Rxy$. In the case of $\exists^{1}X\phi(X)$ we have the usual Peano operator $\iota X\phi(X)$. For the situation $\forall X\exists^{1}Y\phi(X,Y)$ we can denote this by $j(X)=\iota Y\phi(X,Y)$. This corresponds to a genitive: $j(X)$ is read as $X$’s $\phi$ while $j(X)=\epsilon Y\phi(X,Y)$ is read as a $\phi$ of $X$999to simplify the discussion we have overlooked the fact that in Aristotle instead of sentences of the form $\exists X\phi(X)$ and $\forall X\exists Y\phi(X,Y)$ we mainly meet restricted versions such as $\exists X(\psi(X)\&\phi(X))$ and $\forall X\exists Y(\psi(X)\rightarrow(\phi(X,Y))$. For example: ’there is man who knows grammar’ and ’every man has a friend’. The $\epsilon$ construction would then correspond to: ’let $m$ be a man who knows grammar’ and ’let $f$ be a friend of $m$’. See also the discussion in [7] regarding the tis + ekeinos construction for such cases of multiple generality. .
There are not many detailed mathematical proofs in the surviving Aristotelian corpus but the geometrically grounded proofs in the Physics and De Caelo et Mundi (see [9] for an overview of the mathematics of Aristotle) leave little doubt as to the presence of sentences with embedded quantifiers. For instance in Book VII of the Physics101010241a25 in the proof we discussed that everything that moves is moved by something, hapan to kinoumenon hupo tinos anagkê kinesthai, Aristotle makes use of this axiom:
$\exists x(Stops\,xt\rightarrow Stops\,yt)\rightarrow\exists zMoves\,zyt$
where $t$ represents a given time:
In the second place that which is in motion without being moved by anything does not necessarily cease from its motion because something else is at rest, but a thing must be moved by something if the fact of something else having ceased from its motion causes it to be at rest.
As we will discuss in the next section, Aristotle’s theory of definition in the Topics shows traces of a more complex theory involving a theory of relations and in particular there must have been a theory of how definitions of relations were carried out. How would basic relations such as that of $x$ being the paternal grandfather of $y$ ($PGxy$) have been defined ? There is little doubt that Aristotle would have accepted the definition:
This can be written
$PGF\,xy\leftrightarrow\exists z(Father\,xz\enspace\&\enspace Father\,zy)$
Even the most basic reasoning involving such relations would require something like Aristotle’s ekthesis and $\forall^{1}$. This we present as evidence that ekthesis was a fundamental tool of Aristotle’s logic. Let us give an example of one such argument and how Aristotle might have deployed ekthesis in a way strictly analogous to the examples from the Analytics discussed previously. Let us assume the axiom every man has a father $\forall y\exists xFather\,xy$. How would Aristotle have proven that every man has a paternal grandfather, $\forall y\exists xPGF\,xy$ It seems reasonable to suppose that this would have been done by rules for $\forall^{1}$ combined with ekthesis (according to our interpretation) in analogy to the proof 28a24-27 we analysed: by (in this case a repeated use of) a combination of $\exists^{1}$ and $\exists^{1}$. Given a man $m$, we instantiate the axiom to $\exists xFather\,xm$ and then assume via ekthesis that $Fcm$ and then repeat the process for $c$ instead of $m$ to assume $Fc^{\prime}c$ again via ekthesis. We then obtain $Fc^{\prime}c\,\&\,Fcm$ and finish the ektheses by two applications of $\exists^{1}$ to obtain $\exists zPGFzm$. By analogy from the Analytics Aristotle might have expressed this succinctly as follows:
Let $M$ be a man and let $c$ be $M$’s father. Then $c$ has a father and let $c^{\prime}$ be $c$’s father. Then there is some man having $c^{\prime}$ as father and who is the father of $M$, i.e., $c^{\prime}$ is the paternal grandfather of $M$. Thus $M$ has a paternal grandfather.
But we could also reconstruct the argument using the $\epsilon$ interpretation of ekthesis as seen above. In fact since the axiom can be strengthened to every man has a unique father the argument can be expressed very simply using the generalization of the Peano operator considered previously.
We conclude this section by presenting some evidence for the use of quantifier rules over non-monadic predicates and situations embedded quantifiers in Sextus Empiricus. Note that we are evidently not implying that the mere presence of expressions involving multiple generality implies that the ancients had a mechanism for dealing with inferences over such kinds of expression. Rather we are presenting evidence from Sextus of inferences involving multiple generality presented explicitly as inferences111111such inferences, if correct, are no small achievement if one bears in mind how often even the most basic fallacies in propositional logic are committed by people who we would not expect to exhibit faulty reasoning. Arguably natural language competency does necessarily imply logical competency. Showing that multiple generality could have been expressed among the Stoics by a systematic grammatical regimentation as argued in [7] is certainly an important step towards arguing for the historical presence of a Stoic quantifier logic. . It is likely that in late antiquity and the Hellenistic era logic become more syncretic and that there was substantial borrowing between the different schools. Thus even if the passage of Sextus in question seems to primarily concern Stoic logic it is probable that it also reflects indirectly later Peripatetic logic as well in particular with regards to the hypothetical syllogism expounded by Galen. Consider Sextus’ Outlines II, v, 23 (discussing a view of Democritus)[22]:
Democritus says that a human being is what we all know. But as far as this goes, we shall not be acquainted with humans; for we also know dogs — and for that reason dogs will be humans. And there are some humans we do not know — so they will not be humans. Or rather, as far as this conception goes, no-one will be a human; for if Democritus says that a human must be known by everyone, and no human is known by all humans, then no-one will be a human according to him.
We read the deduction of the last sentence as follows: From the premises
$\forall x(Human\,x\leftrightarrow\forall y(Human\,y\rightarrow Know\,yx))$
$\sim\exists x(Human\,x\&\forall y(Human\,y\rightarrow Know\,yx))$
is concluded
$\sim\exists xHuman\,x$
This is a case of multiple generality involving non-monadic predicates. Presumably Sextus is deploying Stoic but maybe also Peripatetic rules to derive the conclusion. If we assume an Aristotelic argumentation then according to our interpretation it might have been carried out as follows. We assume that $\exists xHuman\,x$ then we use a strategy involving $\exists^{1}$. So let $a$ be an existing human,$Human\,a$. We instantiate (1) and through propositional reasoning obtain that
$Human\,a\&\forall y(Human\,y\rightarrow Know(y,a))$
we then apply $\exists^{1}$ to obtain a contradiction $\bot$ from (2). We conclude by applying $\exists^{1}$ to eliminate the hypothesis $Human\,a$ and conclude with $\bot_{c}$. This procedure would seem again quite analogous to the proof by ekthesis in Aristotle we examined.
There also an interesting argument in II, xii, 141:
If some god has told you that this man will be rich, then this man will be rich. But this god (I point, let us suppose, to Zeus) has told you that this man will be rich. Therefore, this man will be rich.[22]
. We interpret this as:
$\forall y((\exists xGodx\&ToldRich\,xy)\rightarrow BeRich\,y)$
$God\,a\&ToldRich\,ab$
therefore
$BeRich\,b$
The natural deduction proof of this inference involves essential use of both $\forall^{1}E$ and $\exists^{1}I$. Here we leave aside the possibility of a more complicated ’intensional’ interpretation of this argument as in [7], that is, a reading of the form ’a god $x$ told $y$ that $[\phi]$’. However [7] does present evidence that the Stoics did indeed consider such complex kinds of lekta.
@@ 7. $\forall x(Ax\rightarrow\sim Bx)$ $\bot_{c}$ 2, 6
Aristotle’s Topics is generally considered to be an early and indeed composite work, reflecting perhaps rather than specifically Aristotelian ideas, common dialectical procedures employed at the Platonic Academy. Indeed the intended purpose of the work is precisely to serve as a debating manual. It is also the part of the Organon that has received less attention from scholars. We hold that such a neglect is undeserved and that the Topics is quite a rich work and of considerable interest in its own right. Indeed it is impossible to do justice to all its aspects in the scope of the present paper. The Topics transcends the strictly ’logical’ concerns of the Analytics as grammatical, semantic, ontological, pragmatic and epistemic aspects equally predominate. From the semantic angle the concern with ambiguity and synonymity is almost an obsession, a fact that is likely explained by preoccupation with facing the rhetorical devices of the sophists. In this section we only consider some aspects of the purely logical content of the Topics. The Topics is a debate manual, but for a very restricted and formalized kind of debate, concerned with set of basic relations among terms [^12] Unfortunately Aristotle’s view of grammar and his theory of terms and concepts is only documented in a few scattered passages (but see chapter 4 of book IV which deals with relative terms and their governed cases.). We know more about Stoic grammar and its grammar-based logic. A basic linguistic entity is the sentence or protasis. A sentence is made up of two linguistic expressions connected by a certain propositional relation. The two expressions are called subject and predicate depending on their position in the propositional relation. Only noun phrases can be a subject but adjectival and verbal constructions can be predicates. These can becomes subjects after a certain syntactic transformation. The original concept of ’verb’ appears to have been that of predicate. A ’term’ oros is what can be a subject. However there is also a sense in which a verb is term to which is added a temporal aspect. The Stoic concept of kategorema seems to fit more exactly the concept of predicate. A predicate can be considered in itself, as unsaturated by any concrete subject. In Plato’s Sophist there is discussion concerning ’sentences’, ’nouns’ and ’verbs’, sentences being combinations of nouns and verbs, and all sentences must have a ’subject’. Plato gives the example ’Theaetetus, with whom I am now speaking, is flying’. This illustrates how Plato’s concept of ’noun’ corresponds to our ’noun phrase’ and very likely to the Aristotelic term or ’oros’ . (which the scholastics called predicabilia), the most important being the species-genus relation, the property relation, the relation of difference (which combined with the species-genus relation yields the further relation of definition) and the relation of accident [^13] It is not clear if this exhausts all the forms of predication. Indeed Aristotle often mentions predication without any further qualification and we are left to wonder if it assumed that this predication must fall within one of the classes above. Also ’accident’ is understood in various different ways.. The second-order predicates $\kappa$ and $\mu$ of the Analytics are a simplification and abstraction of these relations. Following [18] we hold that a topic is a sentence, generally in universally quantified form, expressing properties of these fundamental relations which can be used either to construct or refute a thesis involving one of the predicabilia. We hold that second-order logic is required to express the topics as well as the logical equivalences and inferences implicitly assumed to hold amidst related groups of topics [^14] In 100a Aristotle gives the definition of ’syllogism’. The genus is ’logos’. The difference is that in this logos certain premises being posited something different (heteron) from them occurs (sumbainei) out of necessity. We might think that a syllogism thus corresponds to our modern notion of logical ’rule’ (as in natural deduction or sequent calculi). The problem is that we would then have to reject rules such as $\Gamma,A\vdash A$ for $A$ is not different from the premises. Our position is rather that ’syllogism’ in this context corresponds not to the modern notion of ’rule’ but to a universally quantified second-order sentence as explained further ahead. . In our framework of ASOL the predicabilia are again second-order binary predicates which we denote by $\gamma$ (species-genus), $\iota$ (property), $\delta$ (difference) and $\sigma$(accident). A very important aspect is that they are no longer extensional, contrary to the case for $\kappa$ and $\mu$, that is, our second-order relations cannot be reduced to expressions in which second-order predicates do not occur. And at the same time relations of extension play an important role in many topics: we need to express that the extension of one predicate is strictly contained in another or that they are equal - and this justifies the second-order logic approach because this task can be carried out with the help of the first-order component of the logic. For example if we have $\gamma AB$ then the extension of $A$ is strictly contained in that of $B$, that is, $\forall xAx\rightarrow Bx$ and $\exists xBx\&\sim Ax$, but this last condition is not sufficient (contrary to the case of $\kappa$) to guarantee that $\gamma AB$. Also noteworthy is that the Topics is not restricted to terms corresponding to first-order monadic predicates, there are ’relation’ terms as well, corresponding to binary or higher-order first-order predicate terms. And we can ask what for Aristotle the ’extension’ of such as relation term would be.
@@
The Topics (which has traces of being a compilation of diverse, not always consistent, texts written at different times) consists of eight books. Book IV is dedicated to $\gamma$, book V to $\iota$ and book VI to definition (including a treatment of $\delta$). Important material relating to $\sigma$ is scattered throughout the first two books. Book IV is the most approachable from a formal point of view and seems to also provide a transition to the Analytics. Let us now recapitulate on our fundamental second-order relations and state their corresponding basic axioms relating to extensions. For convenience we will in general write the topics as open formulas where the monadic predicate variables are assumed to be universally quantified. The most common form of a topic is
@@
For the genus-species relation (a second-order binary predicate $\gamma$) we have:
@@
We have the difference relation $\delta AB$ read as ’$A$ is a difference of genus $B$151515 A careful reading of the text (specially Book VI) reveals that Aristotle uses ’difference’ in three different ways: the difference of a species (relative to a genus or a definition), coordinate differences (of the same genus) and differences that ’belong’ to a given genus. We consider the third to be the most fundamental and define the two previous notions of difference in terms of this last one.. For property [^16] The topics of property set the stage for the topics of definition which is the culmination of the Topics. The notions deployed for property are basically the same one as those deployed for definition but is a more straightforward way. There is a much greater laxness in what can count as a property as compared to definition. We might say that property represents of a more archaic form of definition before the requirement of the a regimentation into a (proximate) genus and difference was adopted. The basic idea is that the property of a term $A$ cannot involve $A$ itself (non-circularity), cannot employ less known terms and cannot contain redundancies. $\iota AB$ read ’$B$ is a property of $A$’ we have
[…]
None of these implications are equivalences [^17] There is no doubt that Aristotle conceives of co-extensive terms which are intentionally different (that is, they are essentially different terms, not only syntactically different.). Likewise the terms ’One’ and ’Being’ are described as being predicated of all other terms and yet intentionally distinct..
We could also venture the following axiom for accident [^18] Aristotle distinguishes between separable and inseparable (or essential) accident. But this seems to make sense primarily for a predication of an individual. $\sigma AB$ read ’$B$ is an accident of $A$’:
@@
However the version of accident $\sigma^{\prime}$ in the passage in 120b suggests we need a modal version of ASOL. An accident is simply a predicate which can both be and not be predicated of an individual which we could also read in ASOL as

### v17

@@ 3.2 The Logic of the Prior Analytics
and call the resulting system ESOL. It seems very plausible that AML(see [16]) can be embedded into ESOL and all the corresponding theorems be derived.
@@ 4.1 Basic framework
Aristotle’s Topics is generally considered to be an early and composite work, reflecting perhaps rather than specifically Aristotelian ideas, common dialectical procedures employed at the Platonic Academy. The purpose of the work is to serve as a debating manual. It is the part of the Organon that has received less attention from scholars. But such neglect is undeserved as the Topics is quite a rich work and of considerable interest in its own right and it is impossible to do justice to all its aspects in the scope of the present paper. The Topics transcends the strictly ’logical’ concerns of the Analytics : grammatical, semantic, ontological, pragmatic and epistemic aspects of debate are treated in equal measure. The concern with ambiguity and synonymity amounts almost to an obsession, a fact that is likely explained by the preoccupation with facing the rhetorical devices of the sophists. In this section we confine ourselves primarily to the purely logical content of the Topics. The Topics is a debate manual, but for a very restricted and formalized kind of debate, concerned with a set of basic relations among terms [^6] Unfortunately Aristotle’s view of grammar and his theory of terms and concepts is only documented in a few scattered passages (but see chapter 4 of book IV which deals with relative terms and their governed cases). A basic linguistic entity is the sentence or protasis. A sentence is made up of two linguistic expressions connected by a certain propositional relation. The two expressions are called subject and predicate depending on their position in the propositional relation. Only noun phrases can be a subject but adjectival and verbal constructions can be predicates. These can become subjects after a certain syntactic transformation. The original concept of ’verb’ appears to have been that of predicate. A ’term’ oros is what can be a subject. However there is also a sense in which a verb is a term to which is added a temporal aspect. The Stoic concept of kategorema seems to fit more closely to the concept of predicate. A predicate can be considered in itself, as unsaturated by any concrete subject. In Plato’s Sophist there is a discussion concerning ’sentences’, ’nouns’ and ’verbs’, sentences being combinations of nouns and verbs, and all sentences must have a ’subject’. Plato gives the example ’Theaetetus, with whom I am now speaking, is flying’. This illustrates how Plato’s concept of ’noun’ corresponds to our ’noun phrase’ and very likely to the Aristotelic term or ’oros’. There is very illuminating material regarding terms and the logical syntax of propositions ot be found in Boethius which is discussed in detail in section 9. (which the scholastics called predicabilia), the most important being the species-genus relation, the property relation, the relation of difference (which combined with the species-genus relation yields the further relation of definition) and the relation of accident [^7] It is not clear if this exhausts all the forms of predication. Indeed Aristotle often mentions predication without any further qualification and we are left to wonder if it assumed that this predication must fall within one of the classes above. Also ’accident’ is understood in various different ways. But see the additional comments at the end of this section.. The second-order predicates $\kappa$ and $\mu$ of the Analytics are a simplification and abstraction of these relations. Following [17] and the testimony of Boethius regarding a topic as a ’propositio maxima’ we argue that a topic is a sentence, generally in universally quantified form, expressing properties of these fundamental relations which can be used either to construct or refute a thesis involving one of the predicabilia. We argue that second-order logic is required to express the topics as well as the logical equivalences and inferences implicitly assumed to hold amidst related groups of topics [^8] In 100a Aristotle gives the definition of ’syllogism’. The genus is ’logos’. The difference is that in this logos certain premises being posited something different (heteron) from them occurs (sumbainei) out of necessity. We might think that a syllogism thus corresponds to our modern notion of logical ’rule’ (as in natural deduction or sequent calculi). The problem is that we would then have to reject rules such as $\Gamma,A\vdash A$ for $A$ is not different from the premises. Our position here is rather that ’syllogism’ in this context corresponds not to the modern notion of ’rule’ but to a universally quantified second-order sentence as explained further ahead. . In our framework of ESOL the predicabilia are again $(1,1)$ predicate term symbols which we denote by $\gamma$ (species-genus), $\iota$ (property), $\delta$ (difference) and $\sigma$(accident). But now they are no longer extensional, contrary to the case for $\kappa$ and $\mu$, that is, our relations cannot be reduced to ’first-order’ expressions. However extension plays an important role in many topics: we need to express that the extension of one predicate is strictly contained in another or that they are equal. For example if we have $\gamma AB$ then the extension of $A$ is strictly contained in that of $B$, that is, $\forall xAx\rightarrow Bx$ and $\exists xBx\&\sim Ax$, but this last condition is not sufficient (contrary to the case of $\kappa$) to guarantee that $\gamma AB$. Also noteworthy is that the Topics is not restricted to terms corresponding to monadic predicate term variables, there are ’relation’ terms as well, corresponding to predicate term variables of higher arity. And we can wonder what for Aristotle the ’extension’ of such as relation term would be.
@@
The Topics (which has traces of being a compilation of diverse, not always consistent, texts written at different times) consists of eight books. Book IV is dedicated to $\gamma$, book V to $\iota$ and book VI to definition (including a treatment of $\delta$). Important material relating to $\sigma$ is scattered throughout the first two books. Book IV is the most approachable from a formal point of view and seems to also provide a transition to the Analytics. For convenience we will in general write the topics as open formulas where the monadic predicate term variables are assumed to be universally quantified. The most common form of a topic is
@@
For the genus-species relation $\gamma$ there is evidence that Aristotle posits:
@@
We have the difference relation $\delta AB$ read as ’$A$ is a difference of genus BB [^9] A careful reading of the text (specially Book VI) reveals that Aristotle uses ’difference’ in three different ways: the difference of a species (relative to a genus or a definition), coordinate differences (of the same genus) and differences that ’belong’ to a given genus. We consider the third to be the most fundamental and define the two previous notions of difference in terms of this last one.. For property [^10] The topics of property set the stage for the topics of definition which is the culmination of the Topics. The notions deployed for property are basically the same one as those deployed for definition but in a more straightforward way. There is a much greater laxness in what can count as a property as compared to definition. We might say that property represents of a more archaic form of definition before the requirement of the a regimentation into a (proximate) genus and difference was adopted. The basic idea is that the property of a term $A$ cannot involve $A$ itself (non-circularity), cannot employ less known terms and cannot contain redundancies. $\iota AB$, read ’$B$ is a property of $A$’ , we have
$\iota AB\rightarrow\forall x(Ax\leftrightarrow Bx)\quad\text{(I)}$
None of these implications are equivalences [^11] There is no doubt that Aristotle conceives of co-extensive terms which are intensionally different (that is, they are essentially different terms, not only syntactically different). Likewise the terms ’One’ and ’Being’ are described as being predicated of all other terms and yet intensionally distinct..
We could also venture the following axiom for accident [^12] Aristotle distinguishes between separable and inseparable (or essential) accident. But this seems to make sense primarily for predication of an individual. $\sigma AB$ read ’$B$ is an accident of $A$’:
@@
However the version of accident $\sigma^{\prime}$ in a passage in 120b suggests we need a modal version of ESOL. An accident is simply a predicate which can both be and not be predicated of an individual which we could also read in ESOL as